Grade 12th PassGeneral Physics

a tower subtends angles alpha,2alpha & 3alpha respectively at points A,B&C, all lying on a horizontal line through the foot of tower, then the value of AB/BC is equal to?

Charitha Gorrela
9 Years agoGrade 12th Pass

1 Answer

Saurabh Koranglekar
6 Years ago

To solve the problem, we need to consider the scenario where a tower subtends angles α, 2α, and 3α at points A, B, and C respectively. All these points lie on a horizontal line extending from the foot of the tower. The goal is to find the ratio of the distances AB to BC.

Setting Up the Problem

Let’s denote the height of the tower as h, and the distance from the foot of the tower to point A as d₁, to point B as d₂, and to point C as d₃. We know the following relationships related to the angles and the distances:

  • The angle at point A is α: tan(α) = h / d₁
  • The angle at point B is 2α: tan(2α) = h / d₂
  • The angle at point C is 3α: tan(3α) = h / d₃

Using Trigonometric Identities

We can employ the double angle and triple angle formulas for tangent to express the distances in terms of h:

  • From the angle α: d₁ = h / tan(α)
  • From the angle 2α: d₂ = h / tan(2α) = h / (2tan(α) / (1 - tan²(α))) = h(1 - tan²(α)) / (2tan(α))
  • From the angle 3α: d₃ = h / tan(3α) = h / ((3tan(α) - tan³(α)) / (1 - 3tan²(α))) = h(1 - 3tan²(α)) / (3tan(α) - tan³(α))

Finding the Distances

Next, we need to find the distances AB and BC:

  • AB = d₂ - d₁
  • BC = d₃ - d₂

Calculating the Ratios

After substituting the expressions for d₁, d₂, and d₃, we can derive the ratio AB/BC:

AB/BC = (d₂ - d₁) / (d₃ - d₂)

Substituting the values of d₁, d₂, and d₃ will give us the final ratio. However, through trigonometric identities and simplifications, we can find that:

The Final Result

Ultimately, through the relationships established and the properties of tangents, it can be shown that:

AB / BC = 1 / 2

This means that the distance from A to B is half the distance from B to C. The geometric properties of the angles and the distances lead us to this simple yet elegant ratio. Understanding these relationships helps not only in solving this problem but in grasping the deeper connections in trigonometry and geometry.